# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

### v2

@@ 4 NUMERICAL EXPERIMENTS
To demonstrate the efficacy of our approach, we conduct various experiments. Here, we focus on conditional posterior and coverage properties, but we also provide further ablation studies to better understand our method’s behaviour. We use the aforementioned combination of DPMMs and copula decomposition in all cases. To estimate $\rho$, we always randomly draw $1000$ data points simulated by the PFN. We use the original learning rate approach in the first three sections, but perform dedicated ablations for benchmarks. To this end, we will refer to our method with the learning rate proposed in Fong et al., (2023) as Ours (default), while the adaptations proposed in Section 3.3 based on theoretical considerations and following the discussion about the biasedness of TabPFN are called Ours (type 1) and Ours (type 2).
@@ 4.2 Convergence Speed
We start by investigating the convergence speed of the MP sampling algorithm. For this, we subsample the concrete dataset (Yeh,, 1998) to generate datasets of different sizes. After standardizing the $n\in\{25,100,250,1000\}$ data points, we visually inspect how the conditional posterior mean distribution of $y_{n+k}$ conditional on fixed features $x$ for $B=30$ instances evolves when running our MP approach for $k\in(0,1000]$ steps. Figure 3 summarizes the results for the default type learning rate and confirms that the different trajectories stabilize within the given budget of $N=1000$ sampling steps independent of $n$. This is confirmed in all of our practical applications, where convergence is observed no later than after $N=500$ steps. Results for different learning rate schedules are given in Section B.2.3.
4.3 Conditional Regression Posterior
@@ 4.4 Benchmark
As an empirical evaluation, we compare our method using the previously presented learning rate approaches on real-world data from the UCI repository (Dua and Graff,, 2017). We further compare against a bootstrap approach, which represents the natural comparison method, as it can also determine uncertainty for estimated parameters from PFN without additional parametric assumptions. In this benchmark, we run a conditional quantile regression ($\alpha\in\{0.5,0.9\}$) on the datasets airfoil ($n=1503$, $p=5$), boston ($n=252$, $p=15$), concrete ($n=1030$, $p=8$, Yeh,, 1998), diabetes ($n=442$, $p=10$, Efron et al.,, 2004), energy ($n=768$, $p=8$, Tsanas and Xifara,, 2012), and fish ($n=908$, $p=6$). These datasets provide an ideal benchmark for PFNs, which are still limited in their scalability to larger datasets. We use $n=100$ data points for training and aim to measure the uncertainty of PFN for the predicted quantile on the remaining data points. In addition to our approaches using the martingale posterior with $B=100$, the non-parametric bootstrap is computed. The bootstrap reruns the PFN method for $R=20$ times by drawing samples from the training dataset with replacement, makes quantile predictions, and computes the empirical 90% confidence interval of these predictions. We use both methods to construct a 90% credible/confidence interval (CI) and check the coverage of the predicted quantile given test dataset features. We also compute the length of these intervals to see how conservative our method is in comparison to a bootstrap approach, as well as the time for computation. We repeat the process five times with different splits.
@@ Results
Figure 4 summarizes the results of our benchmark for $\alpha=0.5$, i.e., conditional median regression. The results are similar to those with smaller dataset size and for the quantile regression case (provided in Tables 3 and 4 in Appendix B), with our approach taking a fraction of the bootstrap approach’s time while yielding notably smaller CIs with better coverage. When comparing the different learning rate schemes, we see that in most cases, it is possible to either decrease the interval width (airfoil, boston, concrete, energy, fish) or improve the coverage (boston, concrete, energy, fish), or both, over the approach proposed in Fong et al., (2023).
Runtime
@@ 4.5 Ablation: Dataset Size
This analysis is particularly interesting in light of our discussion in Section 3.5, suggesting that intervals will not yield nominal coverage, even for $n\to\infty$ due to the non-vanishing bias of TabPFN. Empirically, this is validated in Figure 5, showing that neither the learning rate from Fong et al., (2023) nor our suggested alternative (type 1) obtains nominal coverage and even drifts away from the nominal level for increasing $n$.
@@ 5 Discussion
This work proposes an efficient and principled Bayesian uncertainty quantification method for estimates derived from prior fitted networks. Our method thereby provides epistemic uncertainty complementing the aleatoric uncertainty of PFNs. Epistemic uncertainty methods such as deep ensembles (Lakshminarayanan et al.,, 2017) are a potential alternative, but computationally expensive, as they require training multiple models. In the context of pretrained models, this is usually not a feasible option. In contrast, our method only retrains the martingale posterior routine, not the model itself, which is generally much more efficient.
@@
Some limitations remain in the proposed methodology. As shown below, the MP approach cannot fix severe biases in the initial PPD estimate from the PFN. Further, the current implementation does not offer credible sets that are uniform in the feature $x$, because there is no way to forward sample entire prediction curves. Finally, while the methodology may be applied to non-tabular or non-iid settings, these are currently not covered by our theory and experiments. Some potential extensions are discussed in the following.
Posteriors for Other In-Context Learners
Our theory and algorithm can be transferred directly to other in-context predictors that provide a PPD; for example, when using large language models (LLMs) as a prediction tool for numeric value prediction, such as in Falck et al., (2024), or by generating so-called LLM processes (Requeima et al.,, 2024). Both approaches use a text-based prompt to predict future numeric values, thereby reducing the large token space to, e.g., $[0,1]$ (Falck et al.,, 2024) or a 10-dimensional digit space (Requeima et al.,, 2024).
LLM Proof-of-Concept
To substantiate our previous hypothesis, we provide a proof-of-concept by prompting different LLMs to construct a CDF based on $n=25$ using the setup of Falck et al., (2024). The prompt is given in Section B.2.8. While Falck et al., (2024) investigated whether different LLMs possess the martingale property, we are interested in whether our approach can provide uncertainty for these models. To do so, we simulate i.i.d. data from a Gaussian distribution and provide it to the LLM (GPT-4o, GPT-o3-mini-high, DeepSeek R1) including a context that specifies how the data was generated. We then ask the LLM to generate CDF values $F(y_{i})$ for this distribution for a prespecified grid of $y_{i}$ values. Figure 6 shows the initial output CDF of the LLM as a red dashed line. The underlying true CDF is marked with a dashed-dotted green line, indicating that GPT-4o is unable to predict the CDF, while the others do so correctly. We then run $N=1000$ replications of forward loops of our Algorithm 1 to sample values $y_{(n+1):(n+N)}^{(b)}$ and plot the resulting CDF lines for $y_{1:(n+N)}^{(b)},b=1,\ldots,B$ in blue. These lines indicate the uncertainty of the CDF generated by the LLM. The average CDF resulting from these blue curves is plotted in yellow, indicating that all models have the martingale property (although GPT-4o is not correct in its prediction).
While the blue curves also nicely depict the epistemic uncertainty of the LLM for the predicted CDFs, it becomes clear that the provided method—as expected—cannot counteract biases in the LLM (as is the case for GPT 4o). For more advanced LLMs, however, our approach gives uncertainty quantification akin to what is investigated Requeima et al., (2024), but potentially much cheaper.
Posteriors for LLM-Generated Text
PFNs are in-context learners based on the transformer architecture, just as modern LLMs. LLMs are trained to approximate the PPD $p(y|x,\mathcal{D})$ of the next token $y$ given a context $x$ of previous tokens and the corpus $\mathcal{D}$ they were trained on. Given this similarity, the martingale posterior approach could potentially also be adapted to quantify uncertainty in LLM-generated text. There are, however, two major obstacles. First, simple PPD updates, as in (3), are impossible due to the complex relation between the next and the previous tokens. The complexity of such updates can likely only be matched by an equally strong LLM. Second, Proposition 1 and Proposition 2 fail due to the lack of exchangeability in the generated output tokens. Contraction results like Proposition 3, however, rely on only the martingale property and would remain valid, although the interpretation of the distributions $P_{M}$ for $M\to\infty$ is less clear. Investigating how and to what extent martingale posterior ideas transfer to language models is an exciting path for future research.
Supplementary Materials
Appendix A PROOF OF PROPOSITION 3
@@
Next, we apply the uniform-in-time version of the Hoeffding-Azuma inequality given in Corollary 1 (a) of Howard et al., (2020). It holds
@@
Now Corollary 1 (a) of Howard et al., (2020) with $m=C_{\beta}(n+N)^{1-2\beta})$ yields
@@
Now the claim follows since $P_{M}(y)\to P_{\infty,x}(y)$ almost surely (Proposition 2). ∎

### v3

@@ 4 Numerical Experiments
To demonstrate the efficacy of our approach, we conduct various experiments. To estimate $\rho$ for DPMM-based methods, we always randomly draw $1000$ data points simulated by the PFN. If not specified explicitly, we use the default learning rate approach. We will refer to our method with this learning rate as Ours (default), while the adaptations proposed in Section 3.3 based on theoretical considerations and following the discussion about the biasedness of TabPFN are called Ours (type 1) and Ours (type 2). The original approach by Fong et al. (2023) is referred to as Fong et al. The approach that uses purely TabPFN for the forward process is called Pure TabPFN. Code is made available at https://anonymous.4open.science/r/UQ_for_PFNs.
@@ 4.2 Convergence Speed
We start by investigating the convergence speed of the MP sampling algorithm. For this, we subsample the concrete dataset (Yeh, 1998) to generate datasets of different sizes. After standardizing the $n\in\{25,100,250,1000\}$ data points, we visually inspect how the conditional posterior mean distribution of $y_{n+k}$ conditional on fixed features $x$ for $B=30$ instances evolves when running our MP approach for $k\in(0,1000]$ steps. Figure 3 summarizes the results for the default type learning rate and confirms that the different trajectories stabilize within the given budget of $N=1000$ sampling steps independent of $n$. This is confirmed in all of our practical applications, where convergence is observed no later than after $N=500$ steps. Results for different learning rate schedules are given in Section B.3.2.
4.3 Simulation Study
Our approach enables posterior estimation given new features $x$, complementing PFN’s aleatoric uncertainty with epistemic uncertainty. We demonstrate this using a simulation study based on Bayesian additive models. We generate the true underlying feature effects $f_{j}(x_{j})=B(x_{j})^{\top}\theta_{j}$ for feature $x_{j}$ by drawing parameters $\theta_{j}\sim\mathcal{N}(0,I_{20})$ and using those parameters as spline coefficients for a B-spline basis $B_{j}(x_{j})\in\mathbb{R}^{20}$ of degree 3 with 20 basis functions. The data is then generated by the law $Y|x,\theta\sim\mathcal{N}(\sum_{j=1}^{J}f_{j}(x_{j}),\sigma^{2})$ with mean determined by the additive model predictor and fixed variance $\sigma^{2}$. We use 30 features from which $\varphi\in\{1,3,10\}$ have a non-zero signal while the rest act as noise features with no contribution to $Y|x$. At the same time, we again only work with small to moderate amounts of training data $n\in\{30,100,300\}$ and fix $\sigma^{2}=0.5$. Due to the Gaussian likelihood and prior and the linearity of spline coefficients, the resulting posterior is analytically available and can be used to construct credible intervals for each of the functions $f_{j}$ (referred to as Optimal). This allows evaluating the quality of epistemic uncertainty. We use $N=250$ for all methods.
Table 1: Average conditional coverage across all $x$ locations for different methods (columns) for the simulation study using different $(n,\varphi)$ combinations (rows).
$n$ $\varphi$ Optimal Fong et al. Ours Pure TabPFN
30 1 0.91 (1.99) 0.93 (2.57) 0.90 (2.12) 0.83 (2.60)
30 3 0.74 (2.84) 0.65 (3.09) 0.88 (3.52) 0.87 (3.56)
30 10 0.82 (5.57) 0.99 (6.64) 0.94 (5.83) 0.74 (5.91)
100 1 0.92 (1.85) 0.85 (2.65) 0.89 (1.83) 0.82 (1.92)
100 3 0.93 (2.02) 0.96 (3.54) 0.92 (2.77) 0.87 (2.94)
100 10 0.99 (3.17) 0.05 (1.05) 0.99 (6.12) 0.94 (5.83)
300 1 0.91 (1.76) 0.58 (1.54) 0.90 (1.78) 0.84 (1.75)
300 3 0.91 (1.83) 0.92 (3.09) 0.90 (2.16) 0.67 (2.13)
300 10 0.95 (2.06) 0.93 (5.04) 0.97 (5.10) 0.94 (5.06)
@@ Results
The results are depicted in Figure 1 and Table 1 while detailed results breaking down performance into different learning rates and feature locations can be found in Section B.2. While our approach is able to fit the data-generating process well and its uncertainty aligns with the ground truth posterior, initializing Algorithm 1 with the approach proposed Fong et al. (2023) yields a worse fit and coverage for the depicted function. This is expected as, already as a point estimator, Fong’s predictive is far worse than the state-of-the art performance that TabPFN delivers. The pure TabPFN forward process, which also initializes the PPD as in our approach, also shows worse performance and clearly demonstrates that the missing martingale property lets the PPD drift away from the initial estimate when using a pure TabPFN update instead of our proposal in Algorithm 1. Further note that, while all simulation settings, including our method with all three learning rates, can be run within 10 minutes on a conventional CPU laptop, the Pure TabPFN forward process requires more than 15 hours on a GeForce RTX 4070 Ti GPU.
@@ 4.4 Benchmark
As an empirical evaluation, we compare our method using the previously presented learning rate approaches on real-world data from the UCI repository (Dua and Graff, 2017). We further compare against a bootstrap approach, which represents the natural comparison method, as it can also determine uncertainty for estimated parameters from PFN without additional parametric assumptions. In this benchmark, we run a conditional quantile regression ($\alpha\in\{0.5,0.9\}$) on the datasets airfoil ($n=1503$, $p=5$), boston ($n=252$, $p=15$), concrete ($n=1030$, $p=8$, Yeh, 1998), diabetes ($n=442$, $p=10$, Efron et al., 2004), energy ($n=768$, $p=8$, Tsanas and Xifara, 2012), and fish ($n=908$, $p=6$). These datasets provide an ideal benchmark for PFNs, which are still limited in their scalability to larger datasets. We use $n=100$ data points for training and aim to measure the uncertainty of PFN for the predicted quantile on the remaining data points. In addition to our approaches using the martingale posterior with $B=100$, the non-parametric bootstrap is computed. The bootstrap reruns the PFN method for $R=20$ times by drawing samples from the training dataset with replacement, makes quantile predictions, and computes the empirical 90% confidence interval of these predictions. We use both methods to construct a 90% credible/confidence interval (CI) and check the coverage of the predicted quantile given test dataset features. We also compute the length of these intervals to see how conservative our method is in comparison to a bootstrap approach, as well as the time for computation. We repeat the process five times with different splits.
@@ Results
Figure 4 summarizes the results of our benchmark for $\alpha=0.5$, i.e., conditional median regression. The results are similar to those with smaller dataset size and for the quantile regression case (provided in Tables 8 and 9 in Appendix B), with our approach taking a fraction of the bootstrap approach’s time while yielding notably smaller CIs with better coverage. When comparing the different learning rate schemes, we see that in most cases, it is possible to either decrease the interval width (airfoil, boston, concrete, energy, fish) or improve the coverage (boston, concrete, energy, fish), or both, over Fong et al. (2023).
@@ 4.5 Ablation: Dataset Size
This analysis is particularly interesting in light of our discussion in Section 3.6, suggesting that intervals will not yield nominal coverage, even for $n\to\infty$ due to the non-vanishing bias of TabPFN. Empirically, this is validated in Figure 5, showing that neither the learning rate from Fong et al. (2023) nor our suggested alternative (type 1) obtains nominal coverage and even drifts away from the nominal level for increasing $n$.
@@
4.6 Ablation: Initialization and Rollout
Initialization
To empirically validate that the initialization via PFN is the dominant factor in the learned posterior, we conduct another ablation study in which we swap the TabPFN initialization used in Section 4.4 with the PPD learned by the conditional copula regression method proposed in Fong et al. (2023). The learning-rate schedule, the bandwidth, how it is tuned, and the predictive resampling scheme are the same as in our previous experiments.
Rollout
Furthermore, we also investigate pure TabPFN updates of the PPD as suggested in Ng et al. (2025) using forward sampling of training features via Bayesian bootstrap and $N=1000$ forward samples. This ablation allows us to check the extent to which TabPFN’s missing martingale property discussed in Section 3 and the bias underlying the motivation for our proposals in Section 3.6 affect the posterior quality.
Results
The first three columns in Table 2 show the interval widths together with the coverages for our method and the three proposed $\beta$-schedules. While widths are in a similar range as the ones in Figure 4 with TabPFN initialization—which is to be expected as we control for all other tuning parameters—coverage values are far from the nominal level (90%). This is because the initial PPD by Fong et al. (2023) is a much worse estimator of the true conditional distribution than TabPFN, and confirms Corollary 3.1 and the importance of our TabPFN initialization.
The last column in Table 2 demonstrates that a PPD update purely based on TabPFN does not achieve nominal coverages either, even though the intervals are 5 to 10 times wider than the ones produced by our method. This is caused by interval centers drifting away from the initial estimate as a consequence of the lack of a martingale property.
Table 2: Coverage values with interval lengths in brackets across different datasets (rows) and methods.
Default Type 1 Type 2 Pure TabPFN
airfoil 0.18 (0.29) 0.05 (0.07) 0.12 (0.19) 0.27 (0.68)
boston 0.25 (0.23) 0.08 (0.08) 0.19 (0.18) 0.27 (0.51)
concrete 0.16 (0.30) 0.05 (0.08) 0.11 (0.21) 0.34 (1.02)
diabetes 0.34 (0.29) 0.12 (0.10) 0.27 (0.23) 0.65 (0.79)
energy 0.27 (0.12) 0.09 (0.04) 0.20 (0.09) 0.53 (0.89)
fish 0.24 (0.22) 0.06 (0.06) 0.17 (0.15) 0.54 (0.70)
@@ 5 Discussion
This work proposes an efficient and principled Bayesian uncertainty quantification method for estimates derived from prior fitted networks. Our method thereby provides epistemic uncertainty complementing the aleatoric uncertainty of PFNs. Epistemic uncertainty methods such as deep ensembles (Lakshminarayanan et al., 2017) are a potential alternative, but computationally expensive, as they require training multiple models. In the context of pretrained models, this is usually not a feasible option. In contrast, our method only retrains the martingale posterior routine, not the model itself, which is generally much more efficient.
@@
Some limitations remain in the proposed methodology. In particular, the current implementation does not offer credible sets that are uniform in the feature $x$, because there is no way to forward sample entire prediction curves. However, since our method is fast, allowing us to evaluate hundreds of different feature values within minutes, this is less of a problem with our method and more problematic for approaches such as the one from Ng et al. (2025) proposing pure PFN rollouts that can take hours for individual feature conditions. Our proposed methodology may also be applied to non-tabular or non-iid settings, but this is currently not covered by our theory and experiments. In particular, while our theory and algorithm can be transferred directly to other in-context predictors that provide a PPD, such as large language models (LLMs), this is only possible for certain types of prediction tasks, such as those proposed in Falck et al. (2024); Requeima et al. (2024). In this case, a text-based prompt is used to predict future numeric values, thereby reducing the large token space to, e.g., $[0,1]$ (Falck et al., 2024) or a 10-dimensional digit space (Requeima et al., 2024). To substantiate this hypothesis, we provide a proof-of-concept in Appendix D.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A PROOF OF PROPOSITION 3.3
@@
Next, we apply the uniform-in-time version of the Hoeffding-Azuma inequality given in Corollary 1 (a) of Howard et al. (2020). It holds
@@
Now Corollary 1 (a) of Howard et al. (2020) with $m=C_{\beta}(n+N)^{1-2\beta}$ yields
@@
Now the claim follows since $P_{M}(y)\to P_{\infty,x}(y)$ almost surely (Proposition 3.2). ∎